AN AB INITIO STUDY OF THE Ã ^{2}Π STATE AND THE Ã ^{2}Π ← X̃ ^{2}Σ^{+} ELECTRONIC TRANSITION OF MgNC

\underline{TINA ERICA ODAKA}, TSUNEO HIRANO, Department of Chemistry, Faculty of Science, Ochanomizu University, 2-1-1 Otsuka, Bunkyo-ku, Tokyo 112-8610, Japan; PER JENSEN^{a}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, 2-1-1 Bunkyo, Mito 310-8512, Japan.

We have used the RENNER program system^{b} to make a detailed calculation of the rovibronic energies in the first excited electronic state, Ã ^{2}Π, of the MgNC radical. This calculation was based on recent ab initio surfaces (MR-SDCI(+Q)/[TZ3P+f(Mg), aug-cc-pVQZ(N and C)])^{c} for the Renner-degenerate electronic states. Previous calculations of vibronic energies^{c} employing the same ab initio data in conjunction with perturbation expressions^{d} suggested that an observed band belonging to the Ã ^{2}Π ← X̃ ^{2}Σ^{+} electronic transition^{e} should be reassigned. The present work confirms this conclusion which is further substantiated by the rotational structures calculated in the vibronic states, and by Franck-Condon theory predicting relative intensities. Ã ^{2}Π MgNC affords an example of the “classic” Renner effect^{f} involving component electronic states with linear equilibrium geometries. We present detailed analyses of rovibronic wavefunctions aimed at providing further insight into the nature of the Renner interaction.

aSpecial Visiting Professor of the Japanese Ministry of Education, Science and Culture, October-December 2000. Permanent address: FB 9 - Theoretische Chemie, Bergische Universität - Gesamthochschule Wuppertal, D-42097 Wuppertal, Germany.

bSee, for example, P. Jensen, G. Osmann, and P. R. Bunker, in: “Computational Molecular Spectroscopy” (P. Jensen and P. R. Bunker, eds.), Wiley, Chichester, 2000, and references therein.

cT. E. Odaka, T. Taketsugu, T. Hirano, and U. Nagashima, J. Chem. Phys., submitted for publication.

dJ. T. Hougen and J. P. Jesson, J. Chem. Phys. 38, 1524 (1963)

eR. R. Wright and T. A. Miller, J. Mol. Spectrosc. 194, 219 (1999)

fR. Renner, Z. Physik 92, 172 (1934)